# Towards an Information Theoretic Framework of Context-Based Offline Meta-Reinforcement Learning

### v2

@@ Abstract
As a marriage between offline RL and meta-RL, the advent of offline meta-reinforcement learning (OMRL) has shown great promise in enabling RL agents to multi-task and quickly adapt while acquiring knowledge safely. Among which, context-based OMRL (COMRL) as a popular paradigm, aims to learn a universal policy conditioned on effective task representations. In this work, by examining several key milestones in the field of COMRL, we propose to integrate these seemingly independent methodologies into a unified framework. Most importantly, we show that the pre-existing COMRL algorithms are essentially optimizing the same mutual information objective between the task variable $\bm{M}$ and its latent representation $\bm{Z}$ by implementing various approximate bounds. Such theoretical insight offers ample design freedom for novel algorithms. As demonstrations, we propose a supervised and a self-supervised implementation of $I(\bm{Z};\bm{M})$, and empirically show that the corresponding optimization algorithms exhibit remarkable generalization across a broad spectrum of RL benchmarks, context shift scenarios, data qualities and deep learning architectures. This work lays the information theoretic foundation for COMRL methods, leading to a better understanding of task representation learning in the context of reinforcement learning  [^1] The open-source code is at https://github.com/betray12138/UNICORN.git..
@@ 1 Introduction
The ability to swiftly learn and generalize to new tasks is a hallmark of human intelligence. In pursuit of this high level of artificial intelligence (AI), the paradigm of meta-reinforcement learning (RL) proposes to train AI agents in a trial-and-error manner by interacting with multiple external environments. In order to quickly adapt to the unknown, the agents need to integrate prior knowledge with minimal experience (namely the context) collected from the new tasks or environments, without over-fitting to the new data. This meta-RL mechanism has been adopted in many applications such as games [1, 2], robotics [3, 4] and drug discovery [5].
However, for data collection, classical meta-RL usually requires enormous online explorations of the environments [6, 7], which is impractical in many safety-critical scenarios such as healthcare [8] and robotic manipulation [9]. As a remedy, offline RL [10] enables agents to learn from logged experience only, therefore circumventing risky or costly online interactions. Recently, offline meta-RL (OMRL) [11, 12, 13] has emerged as a novel paradigm to significantly extend the applicable range of RL by "killing two birds in one stone": it builds powerful agents that can quickly learn and adapt by meta-learning, while leveraging offline RL mechanism to ensure a safe and efficient optimization procedure. In the context of classical supervised or self-supervised learning, which is de facto offline, OMRL is reminiscent of the multi-task learning [14], meta-training [6] and fine-tuning [15, 16, 17] of pre-trained large models. We envision it as a cornerstone of RL foundation models  [18, 19, 20] in the future.
@@
Despite its effectiveness, FOCAL is reported to be vulnerable to context shifts [21], i.e., when testing on out-of-distribution (OOD) data (Fig. 1). Such problems are particularly challenging for OMRL, since any context shift incurred at test time can not be rectified in the fully offline setting, which may result in severely degraded generalization performance  [21, 22]. To alleviate the problem, follow-up works such as CORRO [23] reformulates the task representation learning of COMRL as maximizing the mutual information $I(\bm{Z};\bm{M})$ between the task variable $\bm{M}$ and its latent $\bm{Z}$. It then approximates $I(\bm{Z};\bm{M})$ by an InfoNCE [24] contrastive loss, where the positive and negative pairs are conditioned on the state-action tuples $(\bm{s},\bm{a})$. Inspired by CORRO, a recently proposed method CSRO [22] introduces an additional mutual information term between $\bm{Z}$ and $(\bm{s},\bm{a})$. By explicitly minimizing it along with the FOCAL objective, CSRO is demonstrated to achieve the state-of-the-art (SOTA) generalization performance on various MuJoCo [25] benchmarks.
@@ 2.1 Preliminaries, Problem Statement and Related Work
However, it is empirically shown that FOCAL struggles to generalize in the presence of context shift  [21]. This problem has been formulated in several earliest studies of OMRL.  Li et al. [28] observed that when there are large divergence of the state-action distributions among the offline datasets, due to shortcut learning [29], the task encoder may learn to ignore the causal information like rewards and spuriously correlate primarily state-action pairs to the task identity, leading to poor generalization performance.  Dorfman et al. [13] concurrently identified a related problem which they termed MDP ambiguity in the context of Bayesian offline RL. An example is illustrated in Figure 1. The problem is further exacerbated in the fully offline setting, as the testing distribution is fixed and cannot be augmented by online exploration, allowing no theoretical guarantee for the context shift.
@@ 2.2 A Unified Information Theoretic Framework
Definition 2.1.
Given an input context variable $\bm{X}\in\mathcal{X}$ and its associated task/MDP random variable $\bm{M}\in\mathcal{M}$, task representation learning in COMRL aims to find a sufficient statistics $\bm{Z}$ of $\bm{X}$ with respect $\bm{M}$.
@@
Definition 2.2.
The dependency graphs of COMRL is given by Fig. 2, where $\bm{X}_{b}$ and $\bm{X}_{t}$ are the behavior-related $(\bm{s},\bm{a})$-component and task-related $(\bm{s}^{\prime},r)$-component of the context $\bm{X}$, with $\bm{X}=(\bm{X}_{t},\bm{X}_{b})$.
@@
For the first graph, $\bm{M}\rightarrow\bm{X}\rightarrow\bm{Z}$ forms a Markov chain, which satisfies $I(\bm{Z};\bm{M}|\bm{X})=0$. To interpret the second graph, given an MDP $M\sim\bm{M}$, the state-action component of $\bm{X}$ is primarily captured by the behavior policy $\pi_{\beta}$: $\bm{s}\sim\mu_{\pi_{\beta}}(\bm{s}),\bm{a}\sim\pi_{\beta}$. The only exception is when tasks differ in initial state distribution $\rho_{0}$ or transition dynamics $T$, in which case the state variable $\bm{S}$ also depends on $\bm{M}$. We therefore define it as the behavior-related component, which should be minimally causally related (dashed lines) to $\bm{M}$ and $\bm{Z}$. Moreover, when $\bm{X}_{b}$ is given, $\bm{X}_{t}$ is fully characterized by the transition function $T:(\bm{s},\bm{a})\rightarrow\bm{s}^{\prime}$ and reward function $R:(\bm{s},\bm{a})\rightarrow r$ of $M$, which should be maximally causally related (solid lines) to $\bm{M}$ and $\bm{Z}$ and therefore be defined as the task-related component. Accordingly, we name $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ and $I(\bm{Z};\bm{X}_{b})$ the primary and lesser causality in our problem respectively. With this prior knowledge and assumption, we present the central theorem of this paper:
Theorem 2.3.
@@
Theorem 2.3 induces several key observations. Firstly, the FOCAL and CORRO objectives operate as an upper bound and a lower bound of the relevant information $I(\bm{Z};\bm{M})$ respectively. Since one would like to maximize $I(\bm{Z};\bm{M})$ according to Definition 2.1, CORRO, which maximizes the lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$, can effectively optimize $I(\bm{Z};\bm{M})$ with theoretical assurance. However, FOCAL which maximizes the upper bound $I(\bm{Z};\bm{X}_{t},\bm{X}_{b})$ provides no guarantee for $I(\bm{Z};\bm{M})$. Mathematically since $I(\bm{Z};\bm{X})=I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})+I(\bm{Z};\bm{X}_{b})$, maximizing the FOCAL objective may instead significantly elevates the lesser causality $I(\bm{Z};\bm{X}_{b})$, which is undesirable since it contains spurious correlation between the task representation $\bm{Z}$ and behavior policy $\pi_{\beta}$. This explains why FOCAL is less robust to context shift compared to CORRO.
Secondly, CSRO as the latest COMRL algorithm, inherently optimizes a linear combination of the FOCAL and CORRO objectives. In the $0\leq\lambda\leq 1$ regime, the CSRO objective becomes a convex interpolation of the upper bound $I(\bm{Z};\bm{X})$ and the lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ of $I(\bm{Z};\bm{M})$, which in essence, enforces a trade-off between the causal ($\bm{Z}$ with $T\,\&\,\rho_{0}$) and spurious ($\bm{Z}$ with $\pi_{\beta}$) correlation contained in $I(\bm{Z};\bm{X}_{b})$. This accounts for the improved performance of CSRO compared to FOCAL and CORRO.
@@ 2.3 Instantiations of UNICORN
Suppose the task label $M$ is given during training and a total of $n_{M}$ training tasks $\{M^{i}\}_{i=1}^{n_{M}}$ are drawn from the task distribution $p(M)$ for meta-training. Under this supervised scenario, by substituting Eqn 8 into 7, we have
@@
Note that $\mathcal{L}_{\textup{UNICORN-SUP}}$ is convex and operates as a finite-sample approximation of $-I(\bm{Z};\bm{M})$, for which we give the following theorem:
@@
Theorem 2.4 bounds the finite-sample estimation error of the empirical risk $\hat{I}(\bm{Z};\bm{M})$ with $n_{M}$ task instances drawn from the real task distribution $p(\bm{M})$. The supervised UNICORN has the merit of directly estimating and optimizing the real objective $I(\bm{Z};\bm{M})$, which requires explicit knowledge of the task label $M^{i}$ and a substantial amount of task instances according to Theorem 2.4. To trade-off computation and performance, we choose to sample 20 training tasks for all RL environments in our experiments.
@@
where $0\leq\alpha\leq 1$ is a hyperparameter. Implementing each term in Eqn 12 allows ample decision choices, such as the contrastive losses in Eqs. 3, 4 and 6 or autoregressive generation via Decision Transformer [33, 34] or RNN [32]. For demonstration, in this paper we employ a contrastive objective $\mathcal{L}_{\textup{FOCAL}}$ as in Eq. 3 for $I(\bm{Z};\bm{X})$ while approximate $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ by reconstruction. By the chain rule of mutual information [35]:
@@
which induces a generative objective $\mathcal{L}_{\textup{recon}}\coloneqq-I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ by reconstructing $\bm{X}_{t}$ with a decoder network $p_{\bm{\theta}}(\cdot|\bm{z},\bm{x_{b}})$ conditioning on $\bm{Z}$ and $\bm{X_{b}}$. As a result, the proposed unsupervised UNICORN objective can be rescaled as Eqn 15. The influence of the hyper-parameter $\frac{\alpha}{1-\alpha}$ is shown in Appendix C.4.
@@
We summarize our learning procedure in Fig. 3 with pseudo-code in Algorithms 1 and 2. A holistic comparison of our proposed algorithms with related contextual meta-RL methods is shown in Table 1. The extra KL divergence can be interpreted as the result of a variational approximation to an information bottleneck [36, 37] that constrains the mutual information between $\bm{Z}$ and $\bm{X}$, which we found unnecessary in our offline setting (see ablation in Table 6.) Behavior regularized actor critic [38] is employed to tackle the bootstrapping error [39] for downstream offline RL implementation.
@@ 3.1 Experimental Setup
We adopt MuJoCo [25] and MetaWorld [3] benchmark to evaluate our method, involving six robotic locomotion tasks which require adaptation across reward functions (HalfCheetah-Dir, HalfCheetah-Vel, Ant-Dir, Reach) or across dynamics (Hopper-Param, Walker-Param). We compare UNICORN with six competitive OMRL algorithms. These baselines are categorized as context-based, gradient-based and transformer-based methods:
Table 2: Average testing returns of UNICORN against baselines on datasets collected by IID and OOD behavior policies. Each result is averaged by 6 random seeds. The best is bolded and the second best is underlined.
@@
MACAW [12] is a gradient-based method, using supervised advantage-weighted regression for both the inner and outer loop of meta-training.
@@
We perform our evaluation in a few-shot manner. For gradient-based methods, we first update the meta-policy with a batch of testing context and then evaluate the agent. For transformer-based methods, we utilize the trajectory segment as the prompt to condition the auto-regressive rollout.

### v3

@@ Abstract
As a marriage between offline RL and meta-RL, the advent of offline meta-reinforcement learning (OMRL) has shown great promise in enabling RL agents to multi-task and quickly adapt while acquiring knowledge safely. Among which, context-based OMRL (COMRL) as a popular paradigm, aims to learn a universal policy conditioned on effective task representations. In this work, by examining several key milestones in the field of COMRL, we propose to integrate these seemingly independent methodologies into a unified framework. Most importantly, we show that the pre-existing COMRL algorithms are essentially optimizing the same mutual information objective between the task variable $\bm{M}$ and its latent representation $\bm{Z}$ by implementing various approximate bounds. Such theoretical insight offers ample design freedom for novel algorithms. As demonstrations, we propose a supervised and a self-supervised implementation of $I(\bm{Z};\bm{M})$, and empirically show that the corresponding optimization algorithms exhibit remarkable generalization across a broad spectrum of RL benchmarks, context shift scenarios, data qualities and deep learning architectures. This work lays the information theoretic foundation for COMRL methods, leading to a better understanding of task representation learning in the context of reinforcement learning. Given its generality, we envision our framework as a promising offline pre-training paradigm of foundation models for decision making. [^1] Source code: https://github.com/betray12138/UNICORN.git.
@@ 1 Introduction
The ability to swiftly learn and generalize to new tasks is a hallmark of human intelligence. In pursuit of this high-level artificial intelligence (AI), the paradigm of meta-reinforcement learning (RL) proposes to train AI agents in a trial-and-error manner by interacting with multiple external environments. In order to quickly adapt to the unknown, the agents need to integrate prior knowledge with minimal experience (namely the context) collected from the new tasks or environments, without over-fitting to the new data. This meta-RL mechanism has been adopted in many applications such as games [1, 2], robotics [3, 4] and drug discovery [5].
However, for data collection, classical meta-RL usually requires enormous online explorations of the environments [6, 7], which is impractical in many safety-critical scenarios such as healthcare [8], autonomous driving [9] and robotic manipulation [10]. As a remedy, offline RL [11] enables agents to learn from logged experience only, thereby circumventing risky or costly online interactions. Recently, offline meta-RL (OMRL) [12, 13, 14] has emerged as a novel paradigm to significantly extend the applicability of RL by "killing two birds in one stone": it builds powerful agents that can quickly learn and adapt by meta-learning, while leveraging offline RL mechanism to ensure a secure and efficient optimization procedure. In the context of classical supervised or self-supervised learning, which is de facto offline, OMRL is reminiscent of the multi-task learning [15], meta-training [6] and fine-tuning [16, 17, 18] of pre-trained large models. We envision it as a cornerstone of RL foundation models  [19, 20, 21] in the future.
@@
Despite its effectiveness, FOCAL is reported to be vulnerable to context shifts [22], i.e., when testing on out-of-distribution (OOD) data (Fig. 1). Such problems are particularly challenging for OMRL, since any context shift incurred at test time can not be rectified in the fully offline setting, which may result in severely degraded generalization performance  [22, 23]. To alleviate the problem, follow-up works such as CORRO [24] reformulates the task representation learning of COMRL as maximizing the mutual information $I(\bm{Z};\bm{M})$ between the task variable $\bm{M}$ and its latent $\bm{Z}$. It then approximates $I(\bm{Z};\bm{M})$ by an InfoNCE [25] contrastive loss, where the positive and negative pairs are conditioned on the same state-action tuples $(\bm{s},\bm{a})$. Inspired by CORRO, a recently proposed method CSRO [23] introduces an additional mutual information term between $\bm{Z}$ and $(\bm{s},\bm{a})$. By explicitly minimizing it along with the FOCAL objective, CSRO is demonstrated to achieve the state-of-the-art (SOTA) generalization performance on various MuJoCo [26] benchmarks.
@@ 2.1 Preliminaries, Problem Statement and Related Work
However, it is empirically shown that FOCAL struggles to generalize in the presence of context shift  [22]. This problem has been formulated in several earliest studies of OMRL.  Li et al. [29] observed that when there are large divergence of the state-action distributions among the offline datasets, due to shortcut learning [30], the task encoder may learn to ignore the causal information like rewards and spuriously correlate primarily state-action pairs to the task identity, leading to poor generalization performance.  Dorfman et al. [14] concurrently identified a related problem which they termed MDP ambiguity in the context of Bayesian offline RL. An example is illustrated in Fig. 1. The problem is further exacerbated in the fully offline setting, as the testing distribution is fixed and cannot be augmented by online exploration, allowing no theoretical guarantee for the context shift.
@@ 2.2 A Unified Information Theoretic Framework
Definition 2.1 (Task Representation Learning).
Given an input context variable $\bm{X}\in\mathcal{X}$ and its associated task/MDP random variable $\bm{M}\in\mathcal{M}$, task representation learning in COMRL aims to find a sufficient statistics $\bm{Z}$ of $\bm{X}$ with respect to $\bm{M}$.
@@
Definition 2.2 (Causal Decomposition).
The dependency graphs of COMRL are given by Fig. 2, where $\bm{X}_{b}$ and $\bm{X}_{t}$ are the behavior-related $(\bm{s},\bm{a})$-component and task-related $(\bm{s}^{\prime},r)$-component of the context $\bm{X}$, with $\bm{X}=(\bm{X}_{t},\bm{X}_{b})$.
@@
For the first graph, $\bm{M}\rightarrow\bm{X}\rightarrow\bm{Z}$ forms a Markov chain, which satisfies $I(\bm{Z};\bm{M}|\bm{X})=0$. To interpret the second graph, given an MDP $M\sim\bm{M}$, the state-action component of $\bm{X}$ is primarily captured by the behavior policy $\pi_{\beta}$: $\bm{s}\sim\mu_{\pi_{\beta}}(\bm{s}),\bm{a}\sim\pi_{\beta}$. The only exception is when tasks differ in initial state distribution $\rho_{0}$ or transition dynamics $T$, in which case the state variable $\bm{S}$ also depends on $\bm{M}$. We therefore define it as the behavior-related component, which should be weakly causally related (dashed lines) to $\bm{M}$ and $\bm{Z}$. Moreover, when $\bm{X}_{b}$ is given, $\bm{X}_{t}$ is fully characterized by the transition function $T:(\bm{s},\bm{a})\rightarrow\bm{s}^{\prime}$ and reward function $R:(\bm{s},\bm{a})\rightarrow r$ of $M$, which should be strongly causally related (solid lines) to $\bm{M}$ and $\bm{Z}$ and therefore be defined as the task-related component. Mathematically, we find that rewriting $\bm{X}=(\bm{X}_{t},\bm{X}_{b})$ induces a causal decomposition of the mutual information $I(\bm{Z};\bm{X})$ by the chain rule [32]:
$I(\bm{Z};\bm{X})=I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})+I(\bm{Z};\bm{X}_{b}).$ (7)
We thereby name $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ and $I(\bm{Z};\bm{X}_{b})$ the primary and lesser causality in our problem respectively. With the setup above, we present the central theorem of this paper:
Theorem 2.3 (Central Theorem).
@@
Theorem 2.3 reveals several key observations. Firstly, the FOCAL and CORRO objectives operate as an upper and a lower bound of $I(\bm{Z};\bm{M})$ respectively. Since one would like to maximize $I(\bm{Z};\bm{M})$ according to Definition 2.1, CORRO, which maximizes the lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$, can effectively optimize $I(\bm{Z};\bm{M})$ with theoretical assurance. However, FOCAL which maximizes the upper bound $I(\bm{Z};\bm{X}_{t},\bm{X}_{b})$ provides no guarantee for $I(\bm{Z};\bm{M})$. By Eq. 7, maximizing the FOCAL objective may instead significantly elevate the lesser causality $I(\bm{Z};\bm{X}_{b})$, which is undesirable since it contains spurious correlation between the task representation $\bm{Z}$ and behavior policy $\pi_{\beta}$. This explains why FOCAL is less robust to context shift compared to CORRO.
Secondly, CSRO as the latest COMRL algorithm among the three, inherently optimizes a linear combination of the FOCAL and CORRO objectives. In the $0\leq\lambda\leq 1$ regime, the CSRO objective becomes a convex interpolation of the upper bound $I(\bm{Z};\bm{X})$ and the lower bound $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ of $I(\bm{Z};\bm{M})$, which in essence, enforces a trade-off between the causal ($\bm{Z}$ with $T$, $\rho_{0}$) and spurious ($\bm{Z}$ with $\pi_{\beta}$) correlation contained in $I(\bm{Z};\bm{X}_{b})$. This accounts for the improved performance of CSRO compared to FOCAL and CORRO.
@@ 2.3 Instantiations of UNICORN
Suppose a total of $n_{M}$ training tasks $\{M^{i}\}_{i=1}^{n_{M}}$ are drawn from the task distribution $p(M)$ with the task label $M$ given for meta-training. Under this supervised scenario, by substituting Eq. 9 into 8, we have
@@
Note that $\mathcal{L}_{\textup{UNICORN-SUP}}$ is convex and operates as a finite-sample approximation of $-I(\bm{Z};\bm{M})$, for which we derive the following bound:
@@
Theorem 2.4 bounds the finite-sample estimation error of the empirical risk $\hat{I}(\bm{Z};\bm{M})$ with $n_{M}$ task instances drawn from the real task distribution $p(\bm{M})$. The supervised UNICORN has the merit of directly estimating and optimizing the real objective $I(\bm{Z};\bm{M})$, which requires explicit knowledge of the task label $M^{i}$ and a substantial amount of task instances according to Theorem 2.4. For better trade-off of computation and performance, we choose to sample 20 training tasks for all RL environments in our experiments.
@@
where $0\leq\alpha\leq 1$ is a hyperparameter. Implementing each term in Eq. 13 allows ample design choices, such as the contrastive losses in Eqs. 3, 4 and 6 or autoregressive generation via Decision Transformer [35, 36] or RNN [34]. For demonstration, in this paper we employ a contrastive objective $\mathcal{L}_{\textup{FOCAL}}$ as in Eq. 3 for estimating $I(\bm{Z};\bm{X})$ while approximate $I(\bm{Z};\bm{X}_{t}|\bm{X}_{b})$ by reconstruction. By the chain rule:
@@
which induces a generative objective $\mathcal{L}_{\textup{recon}}\coloneqq-I(\bm{X}_{t};\bm{Z},\bm{X}_{b})$ by reconstructing $\bm{X}_{t}$ with a decoder network $p_{\bm{\theta}}(\cdot|\bm{z},\bm{x_{b}})$ conditioning on $\bm{Z}$ and $\bm{X_{b}}$. As a result, the proposed unsupervised UNICORN objective can be rescaled as Eq. 16:
@@
The influence of the hyper-parameter $\frac{\alpha}{1-\alpha}$ is shown in Appendix C.4.
We illustrate our learning procedure in Fig. 3 with pseudo-code in Algorithms 1 and 2. A holistic comparison of our proposed algorithms with related contextual meta-RL methods is shown in Table 1. The extra KL divergence in methods like VariBAD and PEARL can be interpreted as the result of a variational approximation to an information bottleneck [37, 38] that constrains the mutual information between $\bm{Z}$ and $\bm{X}$, which we found unnecessary in our offline setting (see ablation in Table 6.) Behavior regularized actor critic [39] is employed to tackle the bootstrapping error [40] for downstream offline RL implementation.
@@ 3.1 Experimental Setup
Data Collection  Following FOCAL, for each task, we train a SAC [43] agent from scratch and collect its replay buffer as the offline dataset. Hence our default data represent a mixed distribution of behavior policies, ranging from random to expert level. Policy checkpoints are also kept for creating various new datasets in Sections 3.3 and 3.4.
We adopt MuJoCo [26] and MetaWorld [3] benchmarks to evaluate our methods, which involve six robotic locomotion environments that require adaptation across reward functions (HalfCheetah-Dir, HalfCheetah-Vel, Ant-Dir, Reach) or across dynamics (Hopper-Param, Walker-Param). We compare UNICORN with six competitive OMRL algorithms, which are categorized as context-based, gradient-based and transformer-based methods:
Table 2: Average testing returns of UNICORN against baselines on datasets collected by IID and OOD behavior policies. Results are averaged by 6 random seeds. The best is bolded and the second best is underlined.
@@
MACAW [13] is a gradient-based method, using supervised advantage-weighted regression for both the inner and outer loops of meta-training.
@@
Our evaluations are performed in a few-shot manner. For gradient-based methods, we first update the meta-policy with a batch of testing context and then evaluate the agent. For transformer-based methods, we utilize trajectory segments as the prompt to condition the auto-regressive rollout.